# A Quasi-local, Functional Analytic Detection Method for Stationary Limit Surfaces of Black Hole Spacetimes

### v2

@@ IV Event Horizons and Black Hole Entropy
Since the stationary limit surfaces of black hole spacetimes with static regions coincide with their event horizons, our functional analytic detection method can be used to locate (slices of) such a particular type of event horizon as shown in detail in Example III.2 in the preceding section for the black hole and cosmological event horizons of the Schwarzschild–de Sitter black hole spacetime. Moreover, it gives rise to the following functional analytic definition of an event horizon in a black hole spacetime with static regions.
@@
=×$\mathfrak{H}=\mathbb{R}\times\bigl{\{}\textnormal{ext}\bigl{(}\textnormal{det}(\sigma_{\mathcal{H}})^{-1}(0)\bigr{)}\bigr{\}}\times(0,\pi)\times[0,2\pi)\,,$
[…]
@@
For every asymptotically flat black hole spacetime with static regions, the largest zeros of the determinantial polynomials of the principal symbols of the Hamiltonians occur at the location of the outer event horizon, that is, $\textnormal{max}\bigl{(}\textnormal{det}(\sigma_{\mathcal{H}})^{-1}(0)\bigr{)}=r^{+}_{\mathfrak{H}}$. Setting the value of the proportionality constant as $\mathscr{C}_{0}=\pi$, the entropy specified in Equation (23) is in accordance with the usual general relativistic black hole entropy expression $S_{\textnormal{BH}}=A/4$, where $A=4\pi(r^{+}_{\mathfrak{H}})^{2}$ is the outer event horizon area. ∎

### v3

@@ IV Event Horizons and Black Hole Entropy
Since the stationary limit surfaces of black hole spacetimes with static regions coincide with their event horizons, our functional analytic detection method can be used to locate (slices of) this particular type of event horizon as shown in detail in Example III.2 in the preceding section for the black hole and cosmological event horizons of the Schwarzschild–de Sitter black hole spacetime. Moreover, it gives rise to the following functional analytic definition of an event horizon in a black hole spacetime with static regions.
@@
$\mathfrak{H}=\mathbb{R}\times\bigl\{\textnormal{ext}\bigl(\textnormal{det}(\sigma_{\mathcal{H}})^{-1}(0)\bigr)\bigr\}\times(0,\pi)\times[0,2\pi)\,$,
where $\textnormal{det}(\sigma_{\mathcal{H}})^{-1}(0)$ denotes the zero sets of the determinants of the principal symbols of the Hamiltonians and $\textnormal{ext}(\,.\,)\in\bigl\{\textnormal{min}(\,.\,),\textnormal{max}(\,.\,)\bigr\}$.
@@
For every asymptotically flat black hole spacetime with static regions, the largest zeros of the determinantial polynomials of the principal symbols of the Hamiltonians occur at the location of the outer event horizon, that is, $\textnormal{max}\bigl(\textnormal{det}(\sigma_{\mathcal{H}})^{-1}(0)\bigr)=r^{+}_{\mathfrak{H}}$. Setting the value of the proportionality constant to $\mathscr{C}_{0}=\pi$, the entropy specified in Equation (23) is in accordance with the usual general relativistic black hole entropy expression $S_{\textnormal{BH}}=A/4$, where $A=4\pi(r^{+}_{\mathfrak{H}})^{2}$ is the outer event horizon area. ∎